OpenStudy (linda):

A= 180 B=60. find the exact value of cos(A-B)

OpenStudy (amistre64):

do you recall the formula for cos(a-b)?

OpenStudy (amistre64):

cos(a-b) = cos(a)cos(b)+sin(a)sin(b)

OpenStudy (amistre64):

cos(180) = -1 sin(180)=0 cos(60) = 1/2 sin(60) = sqrt(3)/2

OpenStudy (amistre64):

the answer is -1/2

OpenStudy (linda):

thank you!